OpenStudy (anonymous):

y varies inversely as twice x. When x = 2, y = 3. Find y when x = 6.

OpenStudy (anonymous):

\[y = \frac{k}{2x}\] Put x = 2 and y = 3 here: \[k = 3 \times 4 = 12\] Now, you have k = 12, x = 6 now find y: \[y = \frac{12}{12}\] Get y from here..

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

Yes you are right..